# Terrain Diffusion: A Diffusion-Based Successor to Perlin Noise in Infinite, Real-Time Terrain Generation

### v1

@@ Abstract.
For decades, procedural worlds have been built on procedural noise functions such as Perlin noise, which are fast and infinite, yet fundamentally limited in realism and large-scale coherence. We introduce Terrain Diffusion, an AI-era successor to Perlin noise that bridges the fidelity of diffusion models with the properties that made procedural noise indispensable: seamless infinite extent, seed-consistency, and constant-time random access. At its core is InfiniteDiffusion, a novel algorithm for infinite generation, enabling seamless, real-time synthesis of boundless landscapes. A hierarchical stack of diffusion models couples planetary context with local detail, while a compact Laplacian encoding stabilizes outputs across Earth-scale dynamic ranges. An open-source infinite-tensor framework supports constant-memory manipulation of unbounded tensors, and few-step consistency distillation enables efficient generation. Together, these components establish diffusion models as a practical foundation for procedural world generation, capable of synthesizing entire planets coherently, controllably, and without limits.
@@
Figure 1. A region of a world generated with Terrain Diffusion. The leftmost panel spans roughly five million square kilometers, with about 2.2 million square kilometers of land area, comparable to the size of the Congo. Red boxes denote the region shown in the next panel, illustrating coherent terrain generation across four orders of magnitude in scale. Zoom for details.: :
@@ 1. Introduction
Terrain Diffusion bridges this divide, introducing a learned successor to procedural noise that preserves all its core properties: seamless infinite extensibility, seed-consistency, and constant-time random access. To do so, we generalize MultiDiffusion (Bar-Tal et al., 2023) for infinite inference, enabling real-time, coherent, procedural generation.
A hierarchical diffusion stack unifies global and local structure: a coarse planetary model establishes continental structure, refined by a higher-resolution model that introduces mountain ranges, valleys, and local relief. A novel elevation encoding further stabilizes training and inference across the full dynamic range of Earth’s terrain. Few-step consistency distillation (Song et al., 2023; Lu and Song, 2025) enables real-time inference, and an open-source infinite-tensor framework enables constant-memory streaming and composable manipulation of infinite tensors.
Together, these components establish the first learned system capable of streaming an entire planet in real time on consumer GPUs. Terrain Diffusion demonstrates that diffusion models can serve as a practical foundation for infinite, seed-consistent worlds that can be explored interactively, in real-time, and without restrictions.
@@ 2. Related Works
Procedural noise. Procedural terrain traditionally relies on procedural noise such as Perlin or Simplex noise (Perlin, 1985, 2002), often combined with fractal Brownian motion (fBm) (Fournier et al., 1982). These methods remain popular for their controllability, speed, seed-consistency, and infinite tileability, but they lack the large-scale structure of real landscapes. They produce repetitive, texture-like patterns and cannot reproduce complex features such as mountain ranges, branching valleys, or volcanic and glacial landforms without extensive post-processing.
Diffusion and consistency models. Denoising diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) generate samples by iteratively refining noise and are widely used for high-fidelity synthesis. Consistency models (Song et al., 2023) approximate denoised diffusion outputs in one or a few steps, and continuous variants (Lu and Song, 2025) achieve throughput competitive with GANs while retaining most of the quality of full diffusion sampling, enabling real-time use cases.
Learned terrain generation. GAN-based terrain models (Goodfellow et al., 2020; Voulgaris et al., 2021; Spick and Walker, 2019; Beckham and Pal, 2017) can generate convincing local relief, but they operate on fixed crops and do not tile, limiting them to bounded worlds. Diffusion-based terrain and Earth imagery synthesis (Hu et al., 2024; Guérin et al., 2017; Borne-Pons et al., 2025) further improves fidelity and control, but also assumes finite canvases and requires relatively significant compute. Sharma et al. (Sharma et al., 2024) achieves $1024{\times}$ super-resolution with strong global coherence, yet still targets bounded images, does not run in real time, and does not generate elevation. Jain et al. (Jain et al., 2023) is closest to our work, offering infinite terrain generation by sampling diffusion-based tiles and blending them with a Perlin-based kernel. Because tiles are generated independently and the kernel has no awareness of broader context, global coherence and large scale structure remain tied to procedural noise rather than the learned model. In contrast, Terrain Diffusion couples all tiles through a shared global context and fuses tiles through a fully learned, context-aware mechanism. Procedural noise is used only for continental-scale structure, where data is sparse and simple enough that more complex alternatives would provide little benefit while reducing user control.
@@ 3.2. From MultiDiffusion to InfiniteDiffusion
which is the MultiDiffusion update with only the windows intersecting $R$ evaluated. In the finite setting, the full image $J_{t}$ can be generated in advance, making each $J_{t}[R_{i}]$ effectively free. In the infinite setting, precomputing $J_{t}$ is impossible, so evaluating $J_{t}[R_{i}]$ requires recursively invoking the same update. A direct implementation would therefore incur exponentially growing compute, motivating an alternative strategy.
Figure 2. A conceptual visualization of InfiniteDiffusion with sliding windows. The user’s query $J_{0}[R]$ induces a deterministic chain of window queries: computing $J_{0}[R]$ requires a region $J_{1}[R_{0}]$, which in turn requires a region $J_{2}[R_{1}]$. Querying $J_{2}[R_{1}]$ is inexpensive since it corresponds directly to Gaussian noise.:
@@ 3.3. Practical Querying of InfiniteDiffusion
There is only one additional challenge in making InfiniteDiffusion practical in real-world applications. Each query $J_{t-1}[R]$ typically requires a region of $J_{t}$ larger than $R$ itself, as pixels near the edge of $R$ are generated by windows that extend beyond $R$. In the sliding-window case, regions grow quadratically in area, yielding an overall $O(T^{3})$ time complexity for final generation. While this prevents the use of longer noise schedules, we find that the method remains both fast and effective even when $T$ is near 1, provided each $y_{i}$ captures sufficient context. In practice, this is achieved by either using a few-step model, such as a consistency model, or simulating one by running several diffusion steps at a time.
@@ 3.4. Properties of InfiniteDiffusion
Formal proofs for all properties are provided in appendix   A
@@ 3.5. An Open Source Infinite Tensor Framework for Unbounded Inference
To support unbounded generation without exceeding memory limits, we introduce the Infinite Tensor framework, a Python library that enables sliding window computation over tensors with infinite dimensions. It allows models to process arbitrarily large images as if they were standard PyTorch tensors while keeping only the visible region in memory. Each operation is performed through a fixed-sized sliding window that dynamically loads and evicts data as sampling progresses, permitting inference on arbitrarily large scenes with constant memory use.
@@ 4. Data
For efficient processing, data is downloaded and stored in contiguous $2048\times 2048$ tiles, each covering equal surface area at approximately 90m resolution. To maintain uniform ground resolution, tiles are stretched in longitude so that each pixel represents roughly the same area regardless of latitude. This equal-area tiling ensures consistent scale across training and allows all models to train on data without distortion. Finally, 80% of the tiles are randomly assigned to the train set, and the remainder are withheld for validation.
All models are trained on random crops drawn from random tiles, with sampling biased so that 99% of tiles contain at least 1% land, as ocean regions are simpler and lower priority. Each crop is randomly flipped and rotated in 90 degree increments to reflect our goal of generating infinite, directionless terrain, similar in spirit to Perlin noise.
@@ 5.1. Signed Square-Root Transform
Terrain tiles vary significantly in elevation range. Under a fixed diffusion noise schedule, this leads to uneven effective SNR: low-relief regions behave as though exposed to stronger noise, while high-relief terrain is affected much less. To reduce this variation, we apply a signed square-root transform $z\mapsto\mathrm{sign}(z)\sqrt{|z|}$. Figure 3 illustrates its effect by comparing tile means and standard deviations before and after the transform. In the raw elevation space, tiles with higher absolute elevations exhibit larger variance. The transform compresses high-relief values and distributes variance across tiles more uniformly. In practice, this allows us to train with a more focused noise distribution and enhances the visibility of small features, especially coastlines.
@@ 5.2. Stabilization Via Laplacian Encodings
Due to normalization, the large dynamic range of Earth elevations make model errors deceptively large in absolute units. Without the signed-sqrt transform, even relatively small errors of $\sigma=0.01$ can correspond to $\pm 25$m noise after denormalization. Applying the transform lessens the issue at lower elevations, but worsens it at higher ones. While this noise is less obvious in images, it can become very apparent in interactive settings, where the noise is potentially much larger than the viewer or the surrounding scenery. To mitigate this, we predict a Laplacian-based representation comprising a low-frequency component, obtained by downsampling and blurring the original image, and a residual/high-frequency component given by subtracting the upsampled low frequency component from the original image.
Residual errors are over $30\times$ smaller in magnitude due to their lower variance. To clean the low-frequency channel after generation, we decode the noisy low- and high-frequency components $(L+H)$ into a provisional heightmap, then blur and downsample it to re-extract a denoised low-frequency $\hat{L}$, with any high-frequency noise redirected to the residual $\hat{H}$. Final synthesis uses $\hat{L}+H$, so low-frequency errors vanish while high-frequency detail is preserved. In practice, $L\approx\hat{L}$ even under strong synthetic noise, confirming that re-extraction cleanly isolates low-frequency structure. This Laplacian denoising step reduces the FID of the untiled core diffusion model (introduced next) from $23.21$ to $9.33$, and the corresponding consistency model from $75.75$ to $12.71$.
@@ 5.3. A Hierarchical Model
Planetary terrain spans several orders of magnitude in scale, from continental structure to meter-level detail, making one-pass generation infeasible. Several previous works (Sharma et al., 2024; Zhang et al., 2025; Lee et al., 2023; Du et al., 2024; Zhou and Tang, 2024) have shown that MultiDiffusion produces incoherent and repetitive results when poorly conditioned, and InfiniteDiffusion does not natively provide a solution to this. We therefore organize generation into a small hierarchy of models operating at progressively finer resolutions. Each stage refines and conditions on the one above, maintaining planetary coherence while producing realistic local detail. All models share a common EDM2 (Karras et al., 2024b) backbone with the modifications proposed in sCM (Lu and Song, 2025). The hierarchy begins with a coarse planetary model, which generates the basic structure of the world from a rough, procedural or user-provided layout. The next stage is the core latent diffusion model, which transforms that structure into realistic 46km tiles in latent space. Finally, a consistency decoder expands these latents into a high-fidelity elevation map. We visualize the coarse-to-fine pipeline in Figure  4.
@@
Additionally, we optionally condition on a target distribution over tile ”beauty” scores. A linear regressor trained on 150 hand-labeled examples and hand-crafted Fourier-based features ($R^{2}\approx 0.8$) assigns each tile a beauty value; during training, we generate random score histograms, sample a random score from it, select an appropriate tile, and feed the target distribution as conditioning, which later allows users to bias synthesis toward more or less visually striking terrain. All conditioning data is provided as a flat tensor, and injected into the model alongside the noise embedding.
@@ 5.4. Real-Time Planetary Scale Synthesis
While the 46 km tiles are realistic in isolation, large-scale coherence requires conditioning on broader planetary context. To provide this, we introduce a compact coarse diffusion model that generates the channels required for conditioning the core diffusion model. The user provides initial maps for these variables using procedural noise or hand-drawn sketches. During inference, these inputs are corrupted with gaussian noise according to the user’s preference on a per-channel basis, and concatenated against the usual diffusion inputs. The model follows the EDM2 design but with no downsampling or upsampling operations. This limits the receptive field by design, preventing the model from drifting toward the massive continental structures present in Earth data and avoiding conflicts where global priors override user guidance while still allowing strong local corrections.
To support real-time synthesis, all diffusion models, except the coarse model, are distilled into continuous-time consistency models (Lu and Song, 2025). To improve fidelity further, we apply the guidance scheme proposed in AutoGuidance (Karras et al., 2024a). Combined, these stages form a continuous generation pipeline, from planetary context to local detail, capable of on-demand, real-time synthesis.
@@ 6. Results
We evaluate Terrain Diffusion on visual fidelity and latency. All experiments use a single NVIDIA RTX 3090 Ti GPU. InfiniteDiffusion uses a stride of 32 in the core model and a separable linear weight window that is 1 at the window center and decreases linearly to a small value $\epsilon$ at the boundary, which reduces FID from $25.83$ with a constant map to $17.87$. The decoder model uses windows of size 512 with strides 384. Batch size is 32 for the core model and 1 for other components.
@@ 6.1. Visual Fidelity
For quality, we report FID-50k for (1) non-tiled diffusion samples, (2) non-tiled consistency samples, and (3) tiled samples generated with InfiniteDiffusion [^1] FID is computed on central 984×984 crops of the validation tiles to ensure adequate surrounding context. InfiniteDiffusion is applied only to the latent space; tiling the decoder would require impractically large context, and the decoder already produces near seamless outputs.. This isolates base fidelity, the effect of consistency distillation, and the effect of tiling. Evaluations use raw elevation values, with images normalized by centering each tile and scaling by the larger of its value range or 255, ensuring that images are not expanded beyond the native precision of the data. All results use two-step generation for both the core consistency model and the consistency decoder. Table 1 summarizes our results.
For context, we also measure the decoder’s standalone rFID at $512\times 512$ resolution. The one step variant obtains an rFID of 3.08, while the two step variant reaches 1.07. The two step model uses a small intermediate noise level of $\sigma=0.065$, indicating that the second step performs only minor refinements that may be unnecessary for speed sensitive applications.
Table 1. FID-50k comparison for central crops generated with InfiniteDiffusion vs. direct generations. Lower is better.
Tiling Model Type FID $\downarrow$
None Diffusion 9.34
None Consistency 12.71
InfiniteDiffusion Consistency 17.87
6.2. Real Latency: Time to First and Second Tile

### v2

@@ Abstract.
[…]
@@
A four panel composite shows progressively closer views of synthetic terrain generated by Terrain Diffusion. The leftmost panel displays a broad archipelago of small continents with a red square marking a region of interest. The second panel zooms into that region and reveals many mountain ranges and coastal plains, again marked by a red square highlighting a smaller area. The third panel zooms further into the selected mountain region, showing an intricate mountain range with various valleys and complex topography. The rightmost panel presents the closest view, focusing on fine scale topographic structure with sharply defined ridgelines, valleys, and erosion patterns. Refer to caption
Figure 1. A region of a world generated with Terrain Diffusion. The leftmost panel spans roughly five million square kilometers, with about 2.2 million square kilometers of land area, comparable to the size of the Congo. Red boxes denote the region shown in the next panel, illustrating coherent terrain generation across four orders of magnitude in scale. Zoom for details.
@@ 1. Introduction
Terrain Diffusion addresses these limitations through InfiniteDiffusion, a generalization of MultiDiffusion (Bar-Tal et al., 2023) for infinite inference. Our framework retains the functional utility of noise while leveraging diffusion models for realism far beyond the reach of procedural noise.
A hierarchical diffusion stack unifies global and local structure: a coarse planetary model establishes continental structure, refined by higher-resolution models that introduce mountain ranges, valleys, and local relief. A novel elevation encoding further stabilizes training and inference across the full dynamic range of Earth’s terrain. Few-step consistency distillation (Song et al., 2023; Lu and Song, 2025) enables rapid inference, and an open-source infinite-tensor framework enables constant-memory streaming and composable manipulation of infinite tensors.
Together, these components establish the first learned system capable of streaming an entire planet in real time on consumer GPUs. Terrain Diffusion demonstrates that diffusion models can serve as a practical foundation for infinite, seed-consistent worlds that can be explored interactively and without restrictions.
@@ 2. Related Works
Table 1. Capability Matrix. Comparison of InfiniteDiffusion against procedural and generative baselines.
Method Infinite Deterministic Speed Quality
Procedural Noise Instant Low
InfinityGAN (Lin et al., 2022) Fast Limited*
Standard Diffusion Slow High
Auto-Regression $O(n)$ High
[…]
*Prone to training instability and mode collapse.
Procedural noise. Procedural terrain traditionally relies on procedural noise such as Perlin or Simplex noise (Perlin, 1985, 2002), often combined with fractal Brownian motion (fBm) (Fournier et al., 1982). These methods remain popular for their controllability, speed, seed-consistency, and infinite extent, but they lack the large-scale structure of real landscapes. They produce repetitive, texture-like patterns and cannot reproduce complex features such as mountain ranges, branching valleys, or volcanic and glacial landforms without extensive post-processing.
Diffusion and consistency models. Denoising diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) generate samples by iteratively refining noise and are widely used for high-fidelity synthesis. Consistency models (Song et al., 2023) approximate denoised diffusion outputs in one or a few steps, and continuous variants (Lu and Song, 2025) achieve throughput competitive with GANs while retaining most of the quality of full diffusion sampling, enabling interactive use cases.
Learned terrain generation. GAN-based terrain models (Goodfellow et al., 2020; Voulgaris et al., 2021; Spick and Walker, 2019; Beckham and Pal, 2017; Argudo et al., 2018) can generate convincing local relief, but they operate on fixed crops and do not tile, limiting them to bounded worlds. Diffusion-based synthesis (Hu et al., 2024; Guérin et al., 2017; Borne-Pons et al., 2025) further improves fidelity and control, but also assumes finite canvases and requires relatively significant compute. Jain et al. (Jain et al., 2023) is closest to our work, offering infinite terrain generation by sampling diffusion-based tiles and blending them with a Perlin-based kernel. Because tiles are generated independently and the kernel has no awareness of broader context, structure remains tied to perlin noise rather than the learned model. In contrast, Terrain Diffusion couples all tiles through a shared global context and fuses tiles through a fully learned, context-aware mechanism. Procedural noise is used only for defining continental layouts, where data is sparse and simple enough that more complex alternatives would provide little benefit while reducing user control.
@@ 3.2. From MultiDiffusion to InfiniteDiffusion
which is the MultiDiffusion update with only the windows intersecting $R$ evaluated. In the finite setting, the full image $J_{t}$ can be generated in advance, making each $J_{t}[R_{i}]$ effectively free. In the infinite setting, precomputing $J_{t}$ is impossible, so evaluating $J_{t}[R_{i}]$ requires recursively invoking the same update. A naive implementation would therefore incur exponentially growing compute. To avoid this, we cache intermediate window computations across queries.
[…]
Figure 2. A conceptual visualization of InfiniteDiffusion with sliding windows. The user’s query $J_{0}[R]$ induces a deterministic chain of window queries: computing $J_{0}[R]$ requires a region $J_{1}[R_{0}]$, which in turn requires a region $J_{2}[R_{1}]$. Querying $J_{2}[R_{1}]$ is inexpensive since it corresponds directly to Gaussian noise.
@@ 3.3. Practical Querying of InfiniteDiffusion
3.4. Tractability via Truncated $T$
There is one final but critical barrier in making InfiniteDiffusion practical. Each query $J_{t-1}[R]$ typically requires a region of $J_{t}$ larger than $R$ itself, as pixels near the edge of $R$ are generated by windows that extend beyond $R$. In the sliding-window case, regions grow quadratically in area, yielding an overall $O(T^{3})$ time complexity for final generation. For traditional MultiDiffusion, which relies on continuously fusing diffusion paths over dozens of steps, this complexity renders infinite generation effectively impossible.
To overcome this, we redefine $\Phi$ as an arbitrary denoising function, such as a few-step Consistency Model or a sequence of standard diffusion steps, rather than a single atomic step. This insight decouples $T$ from the internal diffusion schedule, which may be much longer.
Crucially, we find that our framework retains striking coherence even when generation is aggressively truncated to $T=1$ or $2$. We hypothesize that because global structure is largely determined from hierarchical conditioning, overlapping windows only need to enforce local consistency that is already largely aligned through shared noise. Additionally, because the models are trained on random crops, they learn an approximately translation invariant representation that bases predictions on local patterns, which are shared across overlapping windows, rather than absolute position, further reducing seams. Together, these factors make the method robust even under minimal iterative refinement.
@@ 3.5. Properties of InfiniteDiffusion
Formal proofs for all properties are provided in appendix   B.
@@ 3.6. An Open Source Infinite Tensor Framework for Unbounded Inference
To support unbounded generation without exceeding memory limits, we introduce the Infinite Tensor framework, a Python library that enables sliding window computation over tensors with infinite dimensions. It allows models to process arbitrarily large images as if they were standard PyTorch tensors while keeping only the visible region in memory. Each operation is performed through a fixed-sized sliding window that dynamically loads and evicts data as sampling progresses, permitting inference on arbitrarily large scenes with bounded memory use.
@@ 4. Data
To enable truly global generation, we construct a seamless global elevation dataset merging land topography from the 90m MERIT DEM (Yamazaki et al., 2017) and ocean bathymetry from ETOPO1 (NOAA National Geophysical Data Center, 2009), supplemented by climatic data from WorldClim (Fick and Hijmans, 2017). We process this data into equal-area tiles by dynamically stretching longitude based on latitude; this ensures that pixel sizes represent a consistent physical area, allowing the diffusion model to learn features with minimal polar distortion. The dataset is split into $\times2048$ training tiles, with specific details on coastline smoothing, bathymetric merging, and sampling heuristics provided in Appendix  D.
@@ 5.1. Signed Square-Root Transform
Terrain tiles vary significantly in elevation range. Under a fixed diffusion noise schedule, this leads to uneven effective SNR: low-relief regions behave as though exposed to stronger noise, while high-relief terrain is affected much less. In the raw elevation space, tiles with higher absolute elevations exhibit larger variance. To reduce this variation, we apply a signed square-root transform $z\mapsto\mathrm{sign}(z)\sqrt{|z|}$. The transform compresses high-relief values and distributes variance across tiles more uniformly. Namely, the correlation between the mean and log standard deviation of tiles decreases from $0.66$ to $0.31$. In practice, this allows us to train with a more focused noise distribution and enhances the visibility of small features, especially coastlines. Additional details in Appendix  E.
@@ 5.2. Stabilization Via Laplacian Encodings
Due to normalization, the large dynamic range of Earth elevations make model errors deceptively large in absolute units. Even relatively small errors of $\sigma=0.01$ can correspond to $\pm 25$m noise after denormalization. While this noise is less obvious in images, it can become very apparent in interactive settings, where the noise is potentially much larger than the viewer or the surrounding scenery. To mitigate this, we predict a Laplacian-based representation comprising a low-frequency component, obtained by downsampling and blurring the original image, and a residual/high-frequency component given by subtracting the upsampled low frequency component from the original image.
Residual errors are over $30\times$ smaller in magnitude due to their lower variance. To clean the low-frequency channel after generation, we decode the noisy low- and high-frequency components $(L+H)$ into a provisional heightmap, then blur and downsample it to re-extract a denoised low-frequency $\hat{L}$, with any high-frequency noise redirected to the residual $\hat{H}$. Final synthesis uses $\hat{L}+H$, so low-frequency errors vanish while high-frequency detail is preserved. In practice, $L\approx\hat{L}$ even under strong synthetic noise, confirming that re-extraction cleanly isolates low-frequency structure. This Laplacian denoising step reduces the FID (Heusel et al., 2018) of the untiled core diffusion model (introduced next) from $21.51$ to $8.11$, and the corresponding consistency model, which powers our final pipeline, from $75.15$ to $12.72$.
@@ 5.3. A Hierarchical Model
Planetary terrain spans several orders of magnitude in scale, from continental structure to meter-level detail, making one-pass generation infeasible. Several previous works (Sharma et al., 2024; Zhang et al., 2025; Lee et al., 2023; Du et al., 2024; Zhou and Tang, 2024) have shown that MultiDiffusion produces incoherent and repetitive results when poorly conditioned, and InfiniteDiffusion does not natively provide a solution to this. We therefore organize generation into a small hierarchy of models operating at progressively finer resolutions. Each stage refines and conditions on the one above, maintaining large-scale coherence while producing realistic local detail. All models share a common EDM2 (Karras et al., 2024b) backbone with the modifications proposed in sCM (Lu and Song, 2025). The hierarchy begins with a coarse planetary model, which generates the basic structure of the world from a rough, procedural or user-provided layout. The next stage is the core latent diffusion model (Rombach et al., 2021), which transforms that structure into realistic 46km tiles in latent space. Finally, a consistency decoder expands these latents into a high-fidelity elevation map. We visualize the coarse-to-fine pipeline in Figure  3.
@@ 5.4. Real-Time Planetary Scale Synthesis
While the 46 km tiles are realistic in isolation, large-scale coherence requires conditioning on broader planetary context. To provide this, we introduce a compact coarse diffusion model that generates the channels required for conditioning the core diffusion model. The user provides initial maps for these variables using hand-drawn sketches or procedural noise, which we found works as well as learned methods. During inference, these inputs are corrupted with gaussian noise according to the user’s preference on a per-channel basis, and concatenated against the usual diffusion inputs. The model follows the EDM2 design but with no downsampling or upsampling operations. This limits the receptive field by design, preventing the model from drifting toward the massive continental structures present in Earth data and avoiding conflicts where global priors override user guidance while still allowing strong local corrections.
To enable real-time streaming, all diffusion models, except the coarse model, are distilled into continuous-time consistency models (Lu and Song, 2025). To improve fidelity further, we apply the guidance scheme proposed in AutoGuidance (Karras et al., 2024a). Combined, these stages form a complete generation pipeline, from planetary context to local detail, capable of on-demand, real-time synthesis.
@@ 6. Results
We evaluate Terrain Diffusion on visual fidelity and latency. All experiments use a single NVIDIA RTX 3090 Ti GPU. InfiniteDiffusion uses $T=2$ and a stride of 32 in the core model, and $T=1$ with a batch size of 1 in the coarse and decoder models. We use a separable linear weight window that is 1 at the window center and decreases linearly to a small value $\epsilon$ at the boundary, which reduces FID from $19.32$ with a constant map to $14.78$. The decoder model uses windows of size 512 with strides 384. Batch size is 16 for the core model and 1 for other components.
@@ 6.1. Visual Fidelity
We perform internal ablations to isolate the effects of our proposed architecture. We calculate FID (Heusel et al., 2018) for (1) non-tiled diffusion samples, (2) non-tiled consistency samples, and (3) tiled samples generated with InfiniteDiffusion [^1] FID is computed on central 984×984 crops of the validation tiles to ensure adequate surrounding context. InfiniteDiffusion is applied only to the latent space; tiling the decoder would require impractically large context, and the decoder already produces near seamless outputs.. This isolates base fidelity, the effect of consistency distillation, and the effect of tiling.
We also compare to Perlin blending (Jain et al., 2023), where tiles are blended with perlin noise, and naive tiling, where tiles are simply concatenated without blending. To verify the necessity of fusing diffusion paths during the generative process, we also evaluate ’naive InfiniteDiffusion.’ In this setup, overlapping windows are generated independently as black boxes but with shared noise and conditioning, and their final outputs are linearly blended. Table  2 shows our results.
Table 2. FID-50k for generations with InfiniteDiffusion vs. other methods. Lower is better. The distilled theoretical lower bound is underlined. The best tiled result is bolded.
Tiling Distilled FID $\downarrow$
Perlin Blending (Jain et al., 2023) 196.70
Naive Tiling 74.44
Naive InfiniteDiffusion 35.13
InfiniteDiffusion 14.78
None 12.72
None $\times$ 8.11
Perlin blending yields a high FID of 196.70, confirming that standard procedural blending techniques fail to approximate the statistical distribution of natural topography. The significant fidelity gap between InfiniteDiffusion (FID 14.78) and Naive InfiniteDiffusion (FID 35.13) further demonstrates that large-scale coherence cannot be achieved by merely blending independent patches; it requires fusing intermediate paths iteratively within the latent space. Most notably, InfiniteDiffusion preserves the fidelity of the base consistency model with remarkable accuracy (14.78 vs. 12.72). This marginal FID increase indicates that our windowed fusion strategy, even with T=2, successfully maintains perceptual and spatial continuity, scaling from finite training crops to infinite worlds with little degradation in quality.
For context, we also measure the decoder’s standalone rFID at $512\times 512$ resolution. The one step variant obtains an rFID of $2.83$, while the two step variant reaches $1.07$. Tiling only increases the one-step FID to $2.99$. Evaluations use raw elevation values, with images normalized by centering each tile and scaling by the larger of its value range or 255, ensuring that images are not expanded beyond the native precision of the data. All results use two-step generation for the core consistency model and one-step for the consistency decoder.
6.2. Latency: Time to First and Second Tile